A student has committed the errors which follows the Poisson distribution

with an average rate of 1.5 errors per class.

i) What is the probability that she makes at least 3 errors during one class?

ii) What is the probability that she has committed two ‘error-free’ classes in

the two weeks?


1
Expert's answer
2021-12-21T09:10:23-0500

Let "X" be a random variable representing the errors committed by the student. "X" therefore follows a Poisson distribution with rate "\lambda=1.5" per class.


"i)"

The probability that she makes at least 3 errors during one class is given as,

"p(X\ge 3)=1-p(X\lt3)=1-\{p(X=0)+p(X=1)+p(X=2)\}"

"=1-\{0.2231+0.335+0.2510\}=1-0.8090=0.1912"

Therefore, the probability that she makes at least 3 errors during one class is 0.1912.


"ii)"

In two classes, the rate "\lambda=1.5\times2=3". Thus, the probability that she has committed two ‘error-free’ classes is given by,

"p(X=0)={e^3\times3^0\over0!}=e^3=0.0498"

Therefore, the probability that she has committed two ‘error-free’ classes is 0.0498.